OpenStudy (anonymous):

Write equations for the horizontal and vertical lines passing through the point (-2,7)

OpenStudy (anonymous):

horizontal line= m is 0 using the slope intercept form, y-y1=m(x-x1) y-7=0(x+2) y-7=0 vertical line= m=undefined (or denominator is 0) using the slope intercept form, y-y1=m(x-x1) y-7=(1/0)(x+2) 0=x+2

OpenStudy (tkhunny):

Write equations for the horizontal and vertical lines passing through the point (-2,7) It's horizontal. Why would the y-value change? y = 7. Done! It's vertical. Why would the x-value change? x = -2. Done!

OpenStudy (goformit100):

"A warm Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."